Solve the following system of linear equations by using the inverse matrix
method: (10)
x + y + z = 4
-2x - y + 3z = 1
y + 5z = 9
1
Expert's answer
2020-04-22T14:13:18-0400

Solve the following system of linear equations using an inverse matrix

"x+y+z=4\\
-2x-y+3z=1\\
y+5z=9"

Solution


"A=\begin{pmatrix}
 1 & 1&1 \\
 -2 & -1&3\\
 0&1&5
\end{pmatrix}"

"B=\begin{pmatrix}
 4 \\
 1 \\
 9
\end{pmatrix}"

"X=\begin{pmatrix}
 y \\
 x\\
 z
\end{pmatrix}"

"A\u2219X=B"

that is

X=A-1 ∙B

Find the determinant of matrix A:

"detA=\begin{vmatrix}
 1 & 1&1 \\
 -2 & -1&3\\
 0&1&5

\end{vmatrix}= 1·(-1)·5 + 1·3·0 + 1·(-2)·1 - 1·(-1)·0 - 1·3·1 - 1·(-2)·5 =-5 + 0 - 2 - 0 - 3 + 10 = 0"

Answer: Since the determinant of the matrix is zero, the system has no solution.